Mia needs to make 10 liters of a solution that is 10% salt. Solution A is 5% salt and Solution B is 25% salt. Which system of equations can Mia use to find the number of liters of each solution she should combine? (Let x be the number of liters of Solution A, and y be the number of liters of Solution B.)

Answer by robertb(5830)

0.05x + 0.25(10-x) = 0.10*10,
0.05x + 2.5 - 0.25x = 1,
-0.2x = -1.5,
x = 7.5, # liters of 5% salt solution,
2.5 liters of 25% solution.